# Codes Correcting Two Bursts of Exactly $b$ Deletions

### v2

@@ 2. $(2)$
Suppose that we have found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, where $j_{2}-j_{1}\geq 2$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Given $f^{(0,0)}(\mathbf{x})$, $f^{(0,1)}(\mathbf{x})$, $f^{(1,0)}(\mathbf{x})$ and $f^{(1,1)}(\mathbf{x})$, we can recover $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$ from $\mathbf{x}^{\prime}$.
Proof:
We first consider the case $j_{1}=1$, then $I_{j_{1}}^{L}=\emptyset$ and so we only need to determine $\mathbf{x}_{I_{j_{2}}^{L}}$. Let $s\in\{0,1\}$ be such that $j_{2}\equiv s\pmod{2}$. If $j_{1}\equiv s\pmod{2}$, then by Equation 17, $\mathbf{x}_{I_{j}^{L}}$ is known to us for all $j\in K_{s}\setminus\{j_{2}\}$. By the first line of Equation 18, we have $\mathbf{x}_{I_{j_{2}}^{L}}=\left(f^{(s,0)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{2}\}}\mathbf{x}_{I_{j}^{L}}\right)\pmod{2N_{1}}$ since $0\leq\mathbf{x}_{I_{j_{2}}^{L}}<N_{1}$. If $j_{1}\not\equiv s\pmod{2}$, then $j_{1}+1\equiv s\pmod{2}$. Since $j_{2}-j_{1}\geq 2$, it holds that $j_{1}+1\neq j_{2}$. By Equation 17, $\mathbf{x}_{I_{j}^{L}}$ is known to us for all $j\in K_{s}\setminus\{j_{1}+1,j_{2}\}$. Let $\delta^{(0)}=\left(f^{(s,0)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1}+1,j_{2}\}}\mathbf{x}_{I_{j}^{L}}\right)\pmod{2N_{1}}$ and $\delta^{(1)}=\left(f^{(s,1)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1}+1,j_{2}\}}(j-1)\mathbf{x}_{I_{j}^{L}}\right)\pmod{Q_{1}}$. Then it follows from Equation 18 that
$\mathbf{x}_{I_{j_{1}+1}^{L}}+\mathbf{x}_{I_{j_{2}}^{L}}\equiv\delta^{(0)}\pmod{2N_{1}},$ (19)
$j_{1}\mathbf{x}_{I_{j_{1}+1}^{L}}+(j_{2}-1)\mathbf{x}_{I_{j_{2}}^{L}}\equiv\delta^{(1)}\pmod{Q_{1}}.$ (20)
Since $0\leq\mathbf{x}_{I_{j_{1}+1}^{L}},\mathbf{x}_{I_{j_{2}}^{L}}<N_{1}$ and $0\leq\delta^{(0)}<2N_{1}$, we conclude from Equation 19 that $\mathbf{x}_{I_{j_{1}+1}^{L}}+\mathbf{x}_{I_{j_{2}}^{L}}=\delta^{(0)}$. Combining this with Equation 20, we get
$\mathbf{x}_{I_{j_{1}+1}^{L}}+\mathbf{x}_{I_{j_{2}}^{L}}\equiv\delta^{(0)}\pmod{Q_{1}},$ (21)
$j_{1}\mathbf{x}_{I_{j_{1}+1}^{L}}+(j_{2}-1)\mathbf{x}_{I_{j_{2}}^{L}}\equiv\delta^{(1)}\pmod{Q_{1}}.$
Since $0\leq j_{1}<j_{2}-1<Q_{1}$, we can get the unique solution of (21) in the field $\mathbb{F}_{Q_{1}}$: $\left(\mathbf{x}_{I_{j_{1}+1}^{L}}\pmod{Q_{1}},\mathbf{x}_{I_{j_{2}}^{L}}\pmod{Q_{1}}\right)$. Since $0\leq\mathbf{x}_{I_{j_{1}+1}^{L}},\mathbf{x}_{I_{j_{2}}^{L}}<N_{1}\leq Q_{1}$, it holds that $\mathbf{x}_{I_{j_{1}+1}^{L}}=\mathbf{x}_{I_{j_{1}+1}^{L}}\pmod{Q_{1}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}=\mathbf{x}_{I_{j_{2}}^{L}}\pmod{Q_{1}}$. In particular, we obtained $\mathbf{x}_{I_{j_{2}}^{L}}$.
Next, we consider the case $j_{1}>1$. Let $s_{1},s_{2}\in\{0,1\}$ be such that $j_{1}\equiv s_{1}\pmod{2}$ and $j_{2}\equiv s_{2}\pmod{2}$. If $s_{1}=s_{2}=s$, then $\mathbf{x}_{I_{j}^{L}}$ is known to us for all $j\in K_{s}\setminus\{j_{1},j_{2}\}$. Let $\delta^{(0)}=\left(f^{(s,0)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1},j_{2}\}}\mathbf{x}_{I_{j}^{L}}\right)\pmod{2N_{1}}$ and $\delta^{(1)}=\left(f^{(s,1)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1},j_{2}\}}(j-1)\mathbf{x}_{I_{j}^{L}}\right)\pmod{Q_{1}}$. Then it follows from Equation 18 that
$\mathbf{x}_{I_{j_{1}}^{L}}+\mathbf{x}_{I_{j_{2}}^{L}}\equiv\delta^{(0)}\pmod{2N_{1}},$
$(j_{1}-1)\mathbf{x}_{I_{j_{1}}^{L}}+(j_{2}-1)\mathbf{x}_{I_{j_{2}}^{L}}\equiv\delta^{(1)}\pmod{Q_{1}}.$
Similar to the case $j_{1}=1$, we can uniquely solve $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$.
Now suppose $s_{1}\neq s_{2}$. In this case, we have $j_{1}+1\equiv s_{2}\pmod{2}$. Similar to the case when $j_{1}=1$ and $j_{1}\not\equiv j_{2}\pmod{2}$, we can obtain $\mathbf{x}_{I_{j_{2}}^{L}}$. It remains to show how to obtain $\mathbf{x}_{I_{j_{1}}^{L}}$. It is clear that $j_{2}+1\equiv s_{1}\pmod{2}$. Now similar to the case when $j_{1}=1$ and $j_{1}\not\equiv j_{2}\pmod{2}$, we can obtain $\mathbf{x}_{I_{j_{1}}^{L}}$. ∎
@@
Next, we will apply $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ to decode $\mathbf{x}_{I_{j}^{\prime}}$. Recall that $U(\mathbf{x})_{1}$ is $d$-regular. So by definition, we have $\left|I_{j}^{\prime}\right|=(p_{j}-p_{j-1})b\leq db\log n$. Then by Lemma II.2, $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ can be viewed as an integer in $\left[0,N_{2}-1\right]$, where $N_{2}=2^{\log(db\log n)+O_{q,b}(1)}$. Let $Q_{2}\geq\max\{n/b,N_{2}\}$ be the smallest prime. Define
@@
$g^{(1)}(\mathbf{x})=\sum_{j=1}^{r\left(U(\mathbf{x})_{1}\right)}j\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\pmod{Q_{2}}.$
@@
Since $j_{2}-j_{1}\geq 2$, we have $\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\mathbf{x}^{\prime}_{\left[p_{j_{2}-1}b+1-b,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$. So by Lemma II.2, to recover $\mathbf{x}_{I_{j_{1}}^{\prime}}$ and $\mathbf{x}_{I_{j_{2}}^{\prime}}$, it it sufficient to obtain $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$. Since $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$ are known, it is easy to see that $\mathbf{x}_{I_{j}^{\prime}}$ is known for all $j\neq j_{1},j_{2}$. Let $\delta^{(0)}=\left(g^{(0)}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\right)\pmod{2N_{2}}$ and $\delta^{(1)}=\left(g^{(1)}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}j\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\right)\pmod{Q_{2}}$. Then it follows from Equation 22 that
@@
$j_{1}\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+j_{2}\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(1)}\pmod{Q_{2}}.$ (24)
@@
$\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(0)}\pmod{Q_{2}},$ (25)
$j_{1}\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+j_{2}\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(1)}\pmod{Q_{2}}.$
@@
Suppose $q\geq 2$, $n>2b$ and $b>1$. Let $N_{1},N_{2},N_{3},Q_{1},Q_{2}$, $f^{(s,\ell)}(\mathbf{x})$, $g^{(s)}(\mathbf{x})$ (where $s,\ell\in\{0,1\}$), $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above. For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b_{s,0}<2N_{1}$ ($s=0,1$), $0\leq b_{s,1}<Q_{1}$ ($s=0,1$), $0\leq c_{0}<2N_{2}$, $0\leq c_{1}<Q_{2}$ and $0\leq d_{1},d_{2}<N_{3}$, define the code $\mathcal{C}_{2}$ as
$\mathcal{C}_{2}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }d\text{-regular},\\ \eta(U(\mathbf{x})_{1})=a,\\ f^{(s,\ell)}(\mathbf{x})=b_{s,\ell}\text{ for }s,\ell\in\{0,1\},\\ g^{(0)}(\mathbf{x})=c_{0},g^{(1)}(\mathbf{x})=c_{1},\\ h_{e}(\mathbf{x})=d_{1},h_{o}(\mathbf{x})=d_{2}\end{array}\right\}$
Then $\mathcal{C}_{2}$ is a $2$-$b$-burst-deletion correcting code. Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b_{s,\ell}$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that the redundancy of $\mathcal{C}_{2}$ is at most $7\log n+(14b\left\lceil\log q\right\rceil+11)\log\log n+O_{q,b}(1)$:
@@
Let $q=2$. Then Theorem V.1 gives a binary code correcting two $b$-burst-deletions with redundancy at most $7\log n+(14b+11)\log\log n+O_{b}(1)$. Comparing this with Theorem IV.1, we can see that when $2\leq b\leq 16$, the code in Theorem IV.1 has lower redundancy, while when $b\geq 17$, the code in Theorem V.1 has lower redundancy.
@@ VI Non-binary Two-deletion Correcting Codes
$f_{0}(\mathbf{x})=\sum_{j=1}^{r\left(u(\mathbf{x})\right)}\mathsf{DVT}\left(I_{j}\right)\pmod{2N_{1}},$ (27)
$f_{1}(\mathbf{x})=\sum_{j=1}^{r\left(u(\mathbf{x})\right)}j\mathsf{DVT}\left(I_{j}\right)\pmod{Q}.$
@@ VII Conclusion
In this paper, we present constructions of $q$-ary codes capable of correcting two bursts of exactly $b$ deletions with redundancy at most $7\log n+O(\log\log n)$ for all $b\geq 1$ and $q\geq 2$, which improves the redundancy of codes derived from the syndrome compression technique. In particular, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
@@
The following Lemma .1 is derived by slightly modifying the proof of [43, Proposition 3.3.7]. Before showing its proof, we list some facts.

### v3

@@ 2. $(2)$
Suppose that we have found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, where $j_{2}-j_{1}\geq 2$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Given $f(\mathbf{x})$, we can recover $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$ from $\mathbf{x}^{\prime}$.
@@
Next, we will apply $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ to decode $\mathbf{x}_{I_{j}^{\prime}}$. Recall that $U(\mathbf{x})_{1}$ is $d$-regular. So by definition, we have $\left|I_{j}^{\prime}\right|=(p_{j}-p_{j-1})b\leq db\log n$. Then by Lemma II.2, $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ can be viewed as an integer in $\left[0,N_{2}-1\right]$, where $N_{2}=2^{\log(db\log n)+O_{q,b}(1)}$. Let $Q_{1}\geq\max\{n/b,N_{2}\}$ be the smallest prime. Define
@@
$g^{(1)}(\mathbf{x})=\sum_{j=1}^{r\left(U(\mathbf{x})_{1}\right)}j\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\pmod{Q_{1}}.$
@@
Since $j_{2}-j_{1}\geq 2$, we have $\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\mathbf{x}^{\prime}_{\left[p_{j_{2}-1}b+1-b,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$. So by Lemma II.2, to recover $\mathbf{x}_{I_{j_{1}}^{\prime}}$ and $\mathbf{x}_{I_{j_{2}}^{\prime}}$, it it sufficient to obtain $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$ and $\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$. Since $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$ are known, it is easy to see that $\mathbf{x}_{I_{j}^{\prime}}$ is known for all $j\neq j_{1},j_{2}$. Let $\delta^{(0)}=\left(g^{(0)}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\right)\pmod{2N_{2}}$ and $\delta^{(1)}=\left(g^{(1)}(\mathbf{x})-\sum_{j\neq j_{1},j_{2}}j\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)\right)\pmod{Q_{1}}$. Then it follows from Equation 15 that
@@
$j_{1}\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+j_{2}\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(1)}\pmod{Q_{1}}.$ (17)
@@
$\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(0)}\pmod{Q_{1}},$ (18)
$j_{1}\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)+j_{2}\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)\equiv\delta^{(1)}\pmod{Q_{1}}.$
@@
Suppose $q\geq 2$, $n>2b$ and $b>1$. Let $N_{1},N_{2},N_{3},Q_{1}$, $f(\mathbf{x})$, $g^{(0)}(\mathbf{x})$, $g^{(1)}(\mathbf{x})$, $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above. For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b<N_{1}$, $0\leq c_{0}<2N_{2}$, $0\leq c_{1}<Q_{1}$ and $0\leq d_{1},d_{2}<N_{3}$, define the code $\mathcal{C}_{2}$ as
$\mathcal{C}_{2}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }d\text{-regular},\\ \eta(U(\mathbf{x})_{1})=a,\\ f(\mathbf{x})=b,\\ g^{(0)}(\mathbf{x})=c_{0},g^{(1)}(\mathbf{x})=c_{1},\\ h_{e}(\mathbf{x})=d_{1},h_{o}(\mathbf{x})=d_{2}\end{array}\right\}$
Then $\mathcal{C}_{2}$ is a $2$-$b$-burst-deletion correcting code. Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that the redundancy of $\mathcal{C}_{2}$ is at most $6\log n+(14b\left\lceil\log q\right\rceil+11)\log\log n+O_{q,b}(1)$:
@@
Let $q=2$. Then Theorem V.1 gives a binary code correcting two $b$-burst-deletions with redundancy at most $6\log n+(14b+11)\log\log n+O_{b}(1)$. Comparing this with Theorem IV.1, we can see that when $2\leq b\leq 16$, the code in Theorem IV.1 has lower redundancy, while when $b\geq 17$, the code in Theorem V.1 has lower redundancy.
@@ VI Non-binary Two-deletion Correcting Codes
$f_{0}(\mathbf{x})=\sum_{j=1}^{r\left(u(\mathbf{x})\right)}\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)\pmod{2N_{1}},$ (20)
$f_{1}(\mathbf{x})=\sum_{j=1}^{r\left(u(\mathbf{x})\right)}j\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)\pmod{Q}.$
@@ VII Conclusion
In this paper, we present constructions of $q$-ary codes capable of correcting two bursts of exactly $b$ deletions with redundancy at most $7\log n+O(\log\log n)$ for all $b\geq 2$ and $q\geq 2$, which improves the redundancy of codes derived from the syndrome compression technique. Furthermore, when $b=2$, we show that there is a binary code with redundancy at most $6\log n+O\left(\log\log n\right)$. Inspired by these ideas, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
@@
The following Lemma .1 is derived by slightly modifying the proof of [41, Proposition 3.3.7]. Before showing its proof, we list some facts.